What weight of barium hydroxide is
present in 35.0 ml of 0.200N barium
hydroxide?
1
Expert's answer
2020-09-24T08:12:44-0400

mass of Ba(OH)2 present in 35.0 ml of 0.200M barium

"Mass= molarity\u00d7RMM"

Mass=0.002×171.34=0.343g

"Weight=mass \u00d7 g"

Weight=0.343×10

Weight=3.43N